OpenStudy (anonymous):

√-1 ________ (2 - 4i) - (5 +3i) Simplify the expression

OpenStudy (radar):

\[\sqrt{-1}\over (2-4i)(5+3i)\]Is the expression?

OpenStudy (anonymous):

yes it is

OpenStudy (radar):

Or is it\[\sqrt{-1}\over (2-4i) - (5+3i)\]

OpenStudy (anonymous):

yeah that is correct

OpenStudy (radar):

If it is the first\[7-11i \over 340\]If it is the second:\[-(3i+7)\over 58\]

OpenStudy (anonymous):

how did you get 58? I got 40

OpenStudy (radar):

I got 9-49i^2 or that would be 9 -(49)(-1)=9+49=58 but you got me worried let me recheck.

OpenStudy (anonymous):

lol ok

OpenStudy (radar):

(-3-7i)(-3+7i)=9 -49 i^2=9+49=58

OpenStudy (anonymous):

Oh ok, yeah I looked back at my work and I added wrong. so it would be -7 - 3i / 58

OpenStudy (radar):

don't forget\[i ^{2}=-1\]

OpenStudy (anonymous):

So wait.. is it -7 - 3i /58, because before I put i as -1, it was -3i + 7i^2

OpenStudy (radar):

I changed the two signs by placing the sign to the left and on the viniculum

OpenStudy (radar):

Yes, -3i-7 3i+7 ------- or - ----- 58 58 note the minus sign to the left (middle) of the 2nd term.

OpenStudy (anonymous):

Thank you very much :)

OpenStudy (radar):

u r welcome